# Faster Gradient Methods for Highly-smooth Stochastic Bilevel Optimization

### v1

@@ 4 An $\Omega(\epsilon^{-4})$ Lower Bound
In this section, we prove an $\Omega(\epsilon^{-4})$ lower bound for stochastic bilevel optimization via a reduction to single-level optimization. Our lower bound holds for any randomized algorithms A, which consists of a sequence of measurable mappings $\{\texttt{A}_{t}\}_{t=1}^{T}$ that is defined recursively by
@@
We construct a separable bilevel instance such that the upper-level function $f({\bm{x}},{\bm{y}})\equiv f_{\bm{U}}({\bm{x}})$ and its stochastic gradient align with the hard instance in [3], while the lower-level function is the simple quadratic $g({\bm{x}},y)\equiv g(y)=\mu y^{2}/2$ with deterministic gradients. We defer the concrete construction to Appendix E. For this separable bilevel instance, we can show that for any randomized algorithm defined in Eq. (11) that uses oracles $(F_{\bm{U}},G)$, the progress in $\bm{x}$} can be simulated by another randomized algorithm that only uses $F_{\bm{U}}$, meaning that the single-level lower bound [3] also holds.
@@
There exist numerical constants $c>0$ such that for all $\Delta>0$, $L_{1},L_{2},\cdots,L_{p+1}>0$ and $\epsilon\leq c\sqrt{L_{1}\Delta}$, there exists a distribution over the function class ${\mathcal{F}}^{\text{nc-sc}}(L_{0},L_{2},\cdots,L_{p+1},\mu,\Delta)$ and the stochastic gradient estimators satisfying Assumption 2.1, such that any randomized algorithm A defined as Eq. (11) can not find an $\epsilon$-stationary point of $\varphi({\bm{x}})=f({\bm{x}},{\bm{y}}^{*}({\bm{x}}))$ in less than $\Omega(\Delta L_{1}\sigma^{2}\epsilon^{-4})$ SFO calls.
The analysis is simple using our fully separable construction $f({\bm{x}},y)=f_{\bm{U}}({\bm{x}})$ and $g(y)=\mu y^{2}/2$. But we are a bit surprised that our straightforward construction is not used in prior works such as [13]. Below, we give a detailed discussion on the constructions in other works.
@@
Dagréou et al. [13] proved lower bounds for finite-sum bilevel optimization via a similar reduction to single-level optimization. However, the direct extension of their construction in the fully stochastic setting gives $f({\bm{x}},{\bm{y}})=f_{\bm{U}}({\bm{y}})$ and $g({\bm{x}},{\bm{y}})=({\bm{x}}-{\bm{y}})^{2}$, where the high-order derivatives of $f({\bm{x}},{\bm{y}})$ not ${\mathcal{O}}(1)$-Lipschitz in $\bm{y}$} and thus violates our assumptions. Kwon et al. [29] also proved an $\Omega(\epsilon^{-4})$ lower bound for stochastic bilevel optimization. However, their construction $f({\bm{x}},y)=y$ and $g({\bm{x}},y)=(f_{\bm{U}}({\bm{x}})-y)^{2}$ violate the first-order smoothness of $g({\bm{x}},y)$ in $\bm{x}$} when $y$ is far way from $f_{\bm{U}}({\bm{x}})$. In this work, we use a fully separable construction to avoid all the aforementioned issues in other works.
@@ 5 Experiments
In this section, we conduct numerical experiments to verify our theory. We consider the “learn-to-regularize” problem on the “20 Newsgroup” dataset, which is a very standard benchmark in bilevel optimization [19, 21, 33, 10]. In this task, we aim at learning the optimal regularizer for each parameter of a model. We formulate this task into the following bilevel optimization problem:
$\min_{{\bm{x}}\in{\mathbb{R}}^{p}}\ell^{\rm val}({\bm{y}}),\quad{\rm s.t.}\quad{\bm{y}}\in\arg\min_{{\bm{y}}\in{\mathbb{R}}^{q\times p}}\ell^{\rm tr}({\bm{y}})+\|{\bm{W}}_{{\bm{x}}}{\bm{y}}\|^{2},$
where $\bm{x}$} parameterizes the regularization matrix via ${\bm{W}}_{\bm{x}}={\rm diag}(\exp({\bm{x}}))$, $\bm{y}$} parameterizes a linear model that maps $p=130,107$ features to $q=20$ classes, while $\ell^{\rm val}$ and $\ell^{\rm tr}$ denote the validation and training loss, respectively. The whole dataset contains 18,000 samples. We compare our proposed method F2SA-$p$ with both the previous best fully first-order method F2SA and other Hessian-vector-product-based methods stocBiO [21], MRBO and VRBO [46]. We tune $p$ in $\{1,2,\cdots,5\}$ and find that $p=2$ is the optimal choice. One possible reason is that the instance of both $p=1,2$ only requires solving two lower-level problems at each iteration, but the instance of $p\geq 3$ requires solving more than three lower-level problems and may not be concretely efficient. We regard F2SA-2 as an important instantiation of F2SA-$p$ and present its concrete procedure as well as the comparison to F2SA in Appendix F. For all the algorithms, we search the optimal hyperparameters (including $\eta_{x},\eta_{y},\nu$) in a logarithmic scale with base $10$ and present the experiment results in Figure 1, where we also include a line “w/o Reg” that means the baseline without tuning any regularization. It can be observed that: (1) all the hessian-vector-product-based methods are worse than fully first-order methods; (2) the variance reduction technique in VRBO/MRBO is ineffective and may even harm the performance, which also aligns with the findings in [33]; (3) our method F2SA-2 significantly outperforms all the other algorithms. Our preliminary experiment results on the standard benchmark show the potential of F2SA-2 on large-scale bilevel problems.
Figure 1: Performances of different algorithms when learning the optimal regularization.
@@ 6 Conclusions and Future Works
This paper proposes a class of fully first-order method F2SA-$p$ that achieves the $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ SFO complexity for $p$th-order smooth bilevel problems. Our result generalized the best-known $\tilde{\mathcal{O}}({\epsilon}^{-6})$ result [29, 10] from $p=1$ to any $p\in{\mathbb{N}}_{+}$. We also complement our result with an $\Omega(\epsilon^{-4})$ lower bound to show that our method is near-optimal when $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$. Nevertheless, a gap still exists when $p$ is small, and we still do not know how to fill it even for the basic setting $p=1$. Another possible direction is to extend our theory to structured nonconvex-nonconvex bilevel problems studied by many recent works [30, 9, 8, 22, 43, 42].
@@ Appendix A Notations for Tensors
We follow the notation of tensors used by Kolda and Bader [26]. For two $p$-way tensors ${\mathcal{X}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$ and ${\mathcal{Y}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$, their inner product $z=\langle{\mathcal{X}},{\mathcal{Y}}\rangle$ is defined as
@@
If $p$ is even, we choose the generalized central difference. If $p$ is odd, we choose the generalized forward difference. Our choices underpin the following proof.
Proof of Lemma 3.1.
We analyze the case when $p$ is even or odd separately.
@@
For the coefficients $\{\alpha_{j}\}_{j=-p/2}^{p/2}$, we set
$\alpha_{j}=\alpha_{-j},\quad\forall j=0,1,\cdots,p/2.$
@@
$\frac{1}{\nu}\sum_{j=-p/2}^{j=p/2}\alpha_{j}\psi(j\nu)=\underbrace{2\sum_{j=1}^{p/2}\alpha_{j}\sum_{k=1,3,\cdots}^{p/2-1}\frac{j^{k}\nu^{k-1}}{k!}\psi^{(k)}(0)}_{(*)}+{\mathcal{O}}\left(C\nu^{p}\right).$
To let term (*) be equivalent to $\psi^{\prime}(0)$, we let $\{\alpha_{j}\}_{j=1}^{p/2}$ satisfy the following equations:
$2\sum_{j=1}^{p/2}\alpha_{j}j^{k}={\bm{1}}_{k=1},\quad\forall k=1,3,\cdots,p/2-1,$
which is equivalent to let $\{j\alpha_{j}\}_{j=1}^{p/2}$ satisfy the following linear equation
$\begin{pmatrix}1&1&1&\cdots&1\\ 1^{2}&2^{2}&3^{2}&\cdots&(p/2)^{2}\\ 1^{4}&2^{4}&3^{4}&\cdots&(p/2)^{4}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1^{p/2-2}&2^{p/2-2}&3^{p/2-2}&\cdots&(p/2)^{p/2-2}\end{pmatrix}\begin{pmatrix}\alpha_{1}\\ 2\alpha_{2}\\ 3\alpha_{3}\\ \vdots\\ (p/2)\alpha_{p/2}\end{pmatrix}=\begin{pmatrix}1/2\\ 0\\ 0\\ \vdots\\ 0\end{pmatrix}$.
Since the coefficient matrix in the above linear equation is a Vandermonde matrix, we know this equation has a unique solution, which gives the value of $\{\alpha_{j}\}_{j=1}^{p/2}$.
@@
For the coefficients $\{\beta_{j}\}_{j=0}^{p}$, we first let them satisfy the constraint $\sum_{j=0}^{p}\beta_{j}=0$. Then, summing up Eq. (12) with coefficients $\beta_{j}$ gives
$\frac{1}{\nu}\sum_{j=0}^{j=p}\beta_{j}\psi(j\nu)=\underbrace{\sum_{j=0}^{p}\beta_{j}\sum_{k=1}^{p}\frac{j^{k}\nu^{k-1}}{k!}\psi^{(k)}(0)}_{(*)}+{\mathcal{O}}\left(C\nu^{p}\right).$
To let term (*) be equivalent to $\psi^{\prime}(0)$, we let $\{\beta_{j}\}_{j=0}^{p}$ satisfy the following equations:
$\sum_{j=1}^{p}\beta_{j}j^{k}={\bm{1}}_{k=1},\quad\forall k=1,2,\cdots,p,$
which is equivalent to let $\{j\beta\}_{j=1}^{p}$ satisfy the following linear equation
$\begin{pmatrix}1&1&1&\cdots&1\\ 1&2&3&\cdots&p\\ 1^{2}&2^{2}&3^{2}&\cdots&p^{2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1^{p-1}&2^{p-1}&3^{p-1}&\cdots&p^{p-1}\end{pmatrix}\begin{pmatrix}\beta_{1}\\ 2\beta_{2}\\ 3\beta_{3}\\ \vdots\\ p\beta_{p}\end{pmatrix}=\begin{pmatrix}1\\ 0\\ 0\\ \vdots\\ 0\end{pmatrix}$.
As before, the coefficient matrix in the above linear equation is also a Vandermonde matrix. Therefore, we know this equation has a unique solution, which gives the value of $\{\beta_{j}\}_{j=1}^{p}$ and the coefficient $\beta_{0}$ can be calculated by $\beta_{0}=1-\sum_{j=1}^{p}\beta_{j}$. ∎
@@
Based on Eq. (15), we can prove by induction that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$ for all $q=0,\cdots,p$. The induction base for $q=0,1$ is already proved by Chen et al. [10].
Lemma C.2 (Chen et al. [10, Lemma B.2 and B.5]).
@@
Lemma C.3 (Chen et al. [10, Lemma B.1 and Eq. 18]).
@@
3:  Sample random i.i.d indexes $\{(\xi^{y}_{j},\zeta^{y}_{j})\}_{j=0}^{p}$ and $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
4:  for $j=0,\cdots,p$
5:   ${\bm{y}}_{t}^{j,0}={\bm{y}}_{t}^{j}$
6:   for $k=0,1,\cdots,K-1$
@@
10:  end for
11:  Let $\{\beta_{j}\}_{j=0}^{p}$ be the $p$th-order forward difference coefficients defined in Lemma 3.1.
12:  $\Phi_{t}=\frac{1}{S}\sum_{i=1}^{S}\sum_{j=0}^{p}\beta_{j}\left(jF_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\xi^{x}_{i})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\zeta^{x}_{i})}{\nu}\right)$
13:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
14:end for
@@
Lemma D.1 (Cutkosky and Mehta [11, Lemma 2]).
@@
Lemma D.2 (Kwon et al. [29, Lemma C.1]).
Consider the SGD update ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta F_{t}/\|F_{t}\|$ to optimize a $\mu$-strongly convex function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ with $L$-Lipschitz continuous gradients. Let ${\bm{x}}^{*}=\arg\min_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})$ be the unique minimizer to $f$. Suppose $F_{t}$ is an unbiased estimator to $\nabla f({\bm{x}}_{t})$ with variance bounded by $\sigma^{2}$. Setting $\eta<2/(\mu+L)$, we have
@@
Lemma D.3 (Chen et al. [10, Lemma 4.1]).
@@
Lemma D.4 (Chen et al. [10, Lemma B.6]).

### v2

@@ 4 An $\Omega(\epsilon^{-4})$ Lower Bound
In this section, we prove an $\Omega(\epsilon^{-4})$ lower bound for stochastic bilevel optimization via a reduction to single-level optimization. Our lower bound holds for any randomized algorithms A, which can be defined as a sequence of measurable mappings $\{\texttt{A}_{t}\}_{t=1}^{T}$ that is defined recursively by
@@
We construct a separable bilevel instance such that the upper-level function $f({\bm{x}},{\bm{y}})\equiv f_{\bm{U}}({\bm{x}})$ and its stochastic gradient align with the hard instance in (Arjevani et al., 2023), while the lower-level function is the simple quadratic $g({\bm{x}},y)\equiv g(y)=\mu y^{2}/2$ with deterministic gradients. We defer the concrete construction to Appendix E. For this separable bilevel instance, we can show that for any randomized algorithm defined in Eq. (11) that uses oracles $(F_{\bm{U}},G)$, the progress in $\bm{x}$} can be simulated by another randomized algorithm that only uses $F_{\bm{U}}$, meaning that the single-level lower bound (Arjevani et al., 2023) also holds.
@@
There exist numerical constants $c>0$ such that for all $\Delta>0$, $L_{1},\cdots,L_{p+1}>0$ and $\epsilon\leq c\sqrt{L_{1}\Delta}$, there exists a distribution over the function class ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ and the stochastic gradient estimators satisfying Assumption 2.1, such that any randomized algorithm A defined as Eq. (11) can not find an $\epsilon$-stationary point of $\varphi({\bm{x}})=f({\bm{x}},{\bm{y}}^{*}({\bm{x}}))$ in less than $\Omega(\Delta L_{1}\sigma^{2}\epsilon^{-4})$ SFO calls.
Below, we give a detailed discussion on the constructions in related works.
@@
Dagréou et al. (2024) proved lower bounds for finite-sum bilevel optimization via a similar reduction to single-level optimization. However, the direct extension of their construction in the fully stochastic setting gives $f({\bm{x}},{\bm{y}})=f_{\bm{U}}({\bm{y}})$ and $g({\bm{x}},{\bm{y}})=({\bm{x}}-{\bm{y}})^{2}$, where the high-order derivatives of $f({\bm{x}},{\bm{y}})$ not ${\mathcal{O}}(1)$-Lipschitz in $\bm{y}$} and thus violates our assumptions. Kwon et al. (2024a) also proved an $\Omega(\epsilon^{-4})$ lower bound for stochastic bilevel optimization. However, their construction $f({\bm{x}},y)=y$ and $g({\bm{x}},y)=(f_{\bm{U}}({\bm{x}})-y)^{2}$ violate the first-order smoothness of $g({\bm{x}},y)$ in $\bm{x}$} when $y$ is far way from $f_{\bm{U}}({\bm{x}})$. In this work, we use a fully separable construction to avoid all the aforementioned issues in other works.
@@ 5 Experiments
In this section, we conduct numerical experiments to verify our theory. Following (Grazzi et al., 2020; Ji et al., 2021), we consider the “learn-to-regularize” problem of logistic regression (Example 2.2) on the “20 Newsgroup” dataset, which provably satisfies the highly smooth assumption of any order. The dataset contains 18,000 samples, each sample consists of a feature vector in dimensional $130,107$ vector and a label that takes a value in $\{1,\cdots,20\}$. We compare our proposed method F2SA-$p$ with $p\in\{2,3,5,8,10\}$ with both the previous best fully first-order method F2SA (Kwon et al., 2023; Chen et al., 2025b) and other Hessian-vector-product-based methods stocBiO (Ji et al., 2021), MRBO and VRBO (Yang et al., 2021). We also include a baseline “w/o Reg” that means the training result of SGD without tuning any regularization. For all the algorithms, we search the other hyperparameters (including $\eta_{x},\eta_{y},\nu$) in a logarithmic scale with base $10$. We run the algorithms with $K=10$ iterations in the inner loop, and $T=1000$ iterations in the outer loop, and report the test loss/accuracy v.s. the number of outer-loop iterations $t$ in Figure 1. To demonstrate the potential of our methods on nonsmooth nonconvex problems, we also provide additional experiments on a 5-layer multilayer perceptron (MLP) network with ReLU activation in Appendix G. The codes to reproduce our experiments are available online  [^1] https://github.com/TrueNobility303/F2BA.
Figure 1: Performances of different algorithms on Example $2.$
@@ 6 Conclusions and Future Works
This paper proposes a class of fully first-order method F2SA-$p$ that achieves the $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ SFO complexity for $p$th-order smooth bilevel problems. Our result generalized the best-known $\tilde{\mathcal{O}}({\epsilon}^{-6})$ result (Kwon et al., 2024a; Chen et al., 2025b) from $p=1$ to any $p\in{\mathbb{N}}_{+}$. We also complement our result with an $\Omega(\epsilon^{-4})$ lower bound to show that our method is near-optimal when $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$. Nevertheless, a gap still exists when $p$ is small, and how to fill it even for the basic setting $p=1$ is an open problem. Another open problem is whether our theory can be extended our theory to structured nonconvex-nonconvex bilevel problems studied by many recent works (Kwon et al., 2024b; Chen et al., 2024; 2025a; Jiang et al., 2025; Xiao et al., 2023; Xiao and Chen, 2025). In addition, it will also be interesting to further improve the convergence rate of our methods by combining them with variance-reduction (Fang et al., 2018; Cutkosky and Orabona, 2019) or momentum techniques (Fang et al., 2019; Cutkosky and Mehta, 2020).
Acknowledgments
Lesi Chen thanks Jeongyeol Kwon for helpful discussions and Sanyou Mei for pointing out related references. Jingzhao Zhang is supported by Shanghai Qi Zhi Institute Innovation Program, Tsinghua Dushi Funds, and Xiongan AI Institute.
@@ Appendix A Notations for Tensors
We follow the notation of tensors used by Kolda and Bader (2009). For two $p$-way tensors ${\mathcal{X}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$ and ${\mathcal{Y}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$, their inner product $z=\langle{\mathcal{X}},{\mathcal{Y}}\rangle$ is defined as
@@
Now, we analyze the case when $p$ is even or odd separately.
@@
The estimator we use is known as the $p$th-order central difference, whose coefficients are known (Khan et al., 2003). Let $n=p/2$. We select coefficients $\{\alpha_{j}\}_{j=-n}^{n}$ such that
$\alpha_{j}=-\alpha_{-j},\quad\forall j=0,1,\cdots,n.$
@@
$\frac{1}{\nu}\sum_{j=-n}^{j=n}\alpha_{j}\psi(j\nu)=\underbrace{2\sum_{j=1}^{n}\alpha_{j}\sum_{k=1,3,\cdots}^{n-1}\frac{j^{k}\nu^{k-1}}{k!}\psi^{(k)}(0)}_{(*)}+{\mathcal{O}}\left(C\nu^{p}\right).$
To let term (*) be equivalent to $\psi^{\prime}(0)$, we let $\{\alpha_{j}\}_{j=1}^{n}$ satisfy the following equations:
$2\sum_{j=1}^{n}\alpha_{j}j^{k}={\bm{1}}_{k=1},\quad\forall k=1,3,\cdots,n-1,$
which is equivalent to let $\{j\alpha_{j}\}_{j=1}^{n}$ satisfy the following linear equation
$\begin{pmatrix}1&1&1&\cdots&1\\ 1^{2}&2^{2}&3^{2}&\cdots&n^{2}\\ 1^{4}&2^{4}&3^{4}&\cdots&n^{4}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1^{2(n-1)}&2^{2(n-1)}&3^{2(n-1)}&\cdots&n^{2(n-1)}\end{pmatrix}\begin{pmatrix}\alpha_{1}\\ 2\alpha_{2}\\ 3\alpha_{3}\\ \vdots\\ n\alpha_{n}\end{pmatrix}=\begin{pmatrix}1/2\\ 0\\ 0\\ \vdots\\ 0\end{pmatrix}$.
Now we solve this linear equation to determine the values of $\{\alpha_{j}\}_{j=1}^{n}$. Let $\bm{A}$} be the coefficient matrix of this linear equation, and let ${\bm{A}}_{j}$ be the matrix such that the $j$th column of $\bm{A}$} is replaced by the standard unit vector $(1,0,\cdots,0)^{\top}$. By Cramer’s rule, we have
$2j\alpha_{j}=\frac{\det({\bm{A}}_{j})}{\det({\bm{A}})},\quad j=1,\cdots,n.$
By observation, we can find that both $\bm{A}$} and ${\bm{A}}_{j}$ are Vandermonde matrices. Therefore, we can explicitly calculate both $\det({\bm{A}})$ and $\det({\bm{A}}_{j})$ according to the determinant formula of Vandermonde matrices, which leads to
$2j\alpha_{j}=$ $\frac{(-1)^{j-1}\cdot((j-1)!)^{2}~\cdot(n!)^{2}\cdot j!\cdot(2j)!}{(j!)^{2}\cdot(j-1)!\cdot(n-j)!\cdot(2j-1)!\cdot(n+j)!}=\frac{2(-1)^{j-1}(n!)^{2}}{(n-j)!\cdot(n+j)!}.$
Therefore, we have
$\alpha_{j}=\frac{(-1)^{j-1}(n!)^{2}}{j\cdot(n-j)!\cdot(n+j)!},$
from which it is clear that $|\alpha_{j}|\leq 1/j$.
@@
Instead of using the known $p$th-order forward difference (Khan et al., 2003) for which we find that the coefficients will be exponentially large in $p$, we motivate from the $p$th-order central difference above to obtain a stable estimator by leveraging negative points. Let $n=(p+1)/2$. We select coefficients $\{\alpha_{j}\}_{j=1-n}^{n}$ that satisfy the constraint $\sum_{j=1-n}^{n}\alpha_{j}=0$. Then, summing up Eq. (12) with coefficients $\alpha_{j}$ gives
$\frac{1}{\nu}\sum_{j=1-n}^{j=n}\alpha_{j}\psi(j\nu)=\underbrace{\sum_{j=1-n}^{j=n}\alpha_{j}\sum_{k=1}^{p}\frac{j^{k}\nu^{k-1}}{k!}\psi^{(k)}(0)}_{(*)}+{\mathcal{O}}\left(C\nu^{p}\right).$
To let term (*) be equivalent to $\psi^{\prime}(0)$, we let $\{\alpha_{j}\}_{j=1-n}^{n}$ satisfy the following equations:
$\sum_{j=1-n}^{n}\alpha_{j}j^{k}={\bm{1}}_{k=1},\quad\forall k=1,2,\cdots,p,$
which is equivalent to let $\{\hat{\alpha}_{j}\}_{j=(1-n),j\neq 0}^{n}$ satisfy the following linear equation
$\begin{pmatrix}1&1&\cdots&1&1&\cdots&1\\ 1-n&2-n&\cdots&-1&1&\cdots&n\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ (1-n)^{2n+1}&(2-n)^{2n+1}&\cdots&(-1)^{2n+1}&1^{2n+1}&\cdots&n^{2n+1}\end{pmatrix}\begin{pmatrix}\hat{\alpha}_{1-n}\\ \hat{\alpha}_{2-n}\\ \vdots\\ \hat{\alpha}_{n}\end{pmatrix}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}$,
where we denote $\hat{\alpha}_{j}=j\alpha_{j}$ for $j=1-n,\cdots,-1$ and $1,\cdots,n$. Now we solve this linear equation to determine the values of $\{\hat{\alpha}_{j}\}$. Let $\bm{A}$} be the coefficient matrix of this linear equation, and let ${\bm{A}}_{j}$ be the matrix such that the $j$th column of $\bm{A}$} is replaced by the standard unit vector $(1,0,\cdots,0)^{\top}$. By Cramer’s rule, we have
$\hat{\alpha}_{j}=\frac{\det({\bm{A}}_{j})}{\det({\bm{A}})},\quad j=1,\cdots,n.$
Similar to the case of even $p$, both $\bm{A}$} and ${\bm{A}}_{j}$ are Vandermonde matrices. Therefore, we can explicitly calculate both $\det({\bm{A}})$ and $\det({\bm{A}}_{j})$ according to the determinant formula of Vandermonde matrices. Then, for $j=1,\cdots,n$, we can obtain that
$\alpha_{j}=\frac{\hat{\alpha}_{j}}{j}=\frac{(-1)^{j-1}(j-1)!(n-1)!n!j!}{j\cdot j!(j-1)!(n+j-1)!(n-j)!}=\frac{(-1)^{j-1}(n-1)!n!}{j(n+j-1)!(n-j)!}.$
Similarly, for $j=1,\cdots,n-1$, we can obtain that
$\alpha_{-j}=\frac{\hat{\alpha}_{-j}}{-j}=\frac{(-1)^{j}(n-1)!(j-1)!n!j!}{j\cdot j!(n+j-1)!(j-1)!(n+j)!}=\frac{(-1)^{j}(n-1)!n!}{j(n-j-1)!(n+j)!}.$
Therefore, it is easy to see that $|\alpha_{j}|\leq 1/j$ for $j=1,\cdots,n$, and $|\alpha_{-j}|\leq 1/j$ for $j=1,\cdots,n-1$. Finally, we calculate $\alpha_{0}$ from the constraint $\sum_{j=1-n}^{n}\alpha_{j}=0$, which leads to
$\alpha_{0}=-\underbrace{\sum_{j=1}^{n}\frac{(-1)^{j-1}(n-1)!n!}{j(n+j-1)!(n-j)!}}_{S_{1}}-\underbrace{\sum_{j=1}^{n-1}\frac{(-1)^{j}(n-1)!n!}{j(n-j-1)!(n+j)!}}_{S_{2}}.$ (13)
We claim that $\alpha_{0}=-1/n$ and hence $|\alpha_{0}|\leq 1$. We prove our claim by calculating the values of $S_{1}$ and $S_{2}$ to obtain $\alpha_{0}$. For $S_{1}$, we have
$S_{1}=\sum_{j=1}^{n}(-1)^{j-1}\binom{n}{j}\frac{(n-1)!(j-1)!}{(n+j-1)!}.$
The fraction on the right is the Beta function $B(j,n)$, which can be represented as the integral $B(j,n)=\int_{0}^{1}x^{j-1}(1-x)^{n-1}{\rm d}x$. Therefore,
$S_{1}=$ $\sum_{j=1}^{n}(-1)^{j-1}\binom{n}{j}\int_{0}^{1}x^{j-1}(1-x)^{n-1}{\rm d}x$
$=$ $\int_{0}^{1}(1-x)^{n-1}\left(\sum_{j=1}^{n}(-1)^{j-1}\binom{n}{j}x^{j-1}\right){\rm d}x$
$=$ $\int_{0}^{1}\frac{(1-x)^{n-1}}{x}\left(\sum_{j=1}^{n}(-1)^{j-1}\binom{n}{j}x^{j}\right){\rm d}x.$
Substituting the binomial expansion $(1-x)^{n}=1+\sum_{j=0}^{n}\binom{n}{j}(-x)^{j}$, we then have
$S_{1}=\int_{0}^{1}\frac{(1-x)^{n-1}}{x}\left(1-(1-x)^{n}\right){\rm d}x.$
Let y=1xy=1-x. We then have
$S_{1}=\int_{0}^{1}\frac{y^{n-1}}{1-y}(1-y^{n}){\rm d}y.$
Substituting the geometric series sum $\frac{1-y^{n}}{1-y}=\sum_{k=0}^{n-1}y^{k}$, we then have
$S_{1}=\int_{0}^{1}y^{n-1}\left(\sum_{k=0}^{n-1}y^{k}\right){\rm d}y=\int_{0}^{1}\sum_{k=0}^{n-1}y^{n+k-1}{\rm d}y=\sum_{k=0}^{n-1}\int_{0}^{1}y^{n+k-1}{\rm d}y=\sum_{k=0}^{n-1}\frac{1}{n+k}.$
Following similar steps, we can also obtain that
$S_{2}=-\sum_{k=0}^{n-2}\frac{1}{n+k+1}.$
Now, for $\alpha_{0}=-S_{1}-S_{2}$, the summation terms cancel out perfectly, which leads to $\alpha_{0}=-1/n$.
@@
Based on Eq. (16), we can prove by induction that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$ for all $q=0,\cdots,p$. The induction base for $q=0,1$ is already proved by Chen et al. (2025b).
Lemma C.2 (Chen et al. (2025b, Lemma B.2 and B.5)).
@@
Lemma C.3 (Chen et al. (2025b, Lemma B.1 and Eq. 18)).
@@
3:  parallel for $j=-(p-1)/2,\cdots,(p+1)/2$
4:   ${\bm{y}}_{t}^{j,0}={\bm{y}}_{t}^{j}$
5:   for $k=0,1,\cdots,K-1$
6:    Sample random i.i.d indexes $\{(\xi^{y}_{j},\zeta^{y}_{j})\}$.
@@
10:  end parallel for
11:  Sample random i.i.d indexes $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
12:  Let $\{\alpha_{j}\}_{j=-(p-1)/2}^{(p+1)/2}$ be the $p$th-order finite difference coefficients defined in Lemma 3.1.
13:  $\Phi_{t}=\frac{1}{S}\sum_{i=1}^{S}\sum_{j=-(p-1)/2}^{(p+1)/2}\alpha_{j}\left(jF_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\xi^{x}_{i})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\zeta^{x}_{i})}{\nu}\right)$
14:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
15:end for
@@
Lemma D.1 (Cutkosky and Mehta (2020, Lemma 2)).
@@
Lemma D.2 (Kwon et al. (2024a, Lemma C.1)).
Consider the SGD update ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta F_{t}$ to optimize a $\mu$-strongly convex function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ with $L$-Lipschitz continuous gradients. Let ${\bm{x}}^{*}=\arg\min_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})$ be the unique minimizer to $f$. Suppose $F_{t}$ is an unbiased estimator to $\nabla f({\bm{x}}_{t})$ with variance bounded by $\sigma^{2}$. Setting $\eta<2/(\mu+L)$, we have
@@
Lemma D.3 (Chen et al. (2025b, Lemma 4.1)).
@@
Lemma D.4 (Chen et al. (2025b, Lemma B.6)).